OpenStudy (anonymous):

Solve the system of equations(use either substitution or elimination). SHOW ALL WORK 3x-2y=10 2y=6x-4

OpenStudy (anonymous):

I just did a bunch of these.

OpenStudy (anonymous):

I am doing it for extra credit

OpenStudy (anonymous):

ok ill medal you f u help me out

OpenStudy (anonymous):

Give me a minute and I might be able to give answer. I am working it out on paper.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

also I need the step by step process because it says to show all work

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Okay so let me type all this out.

OpenStudy (anonymous):

\[3x-2y=10\]\[2y=6x-4\]

OpenStudy (anonymous):

So to find the answer, I am using the substitution method. It is easier for me.

OpenStudy (anonymous):

So to start off, you need to move your variables. So, (give me a sec)

OpenStudy (anonymous):

\[3x-2y=10\] and \[6x-2y=-4\]

OpenStudy (anonymous):

What do you think we do next?

OpenStudy (anonymous):

Idk thts why I am asking you for help

OpenStudy (anonymous):

Okay. I was asking because you need to know this. If you keep asking in here istead of knowing, you won't be prepared for eos exams. Give me a minute and I will type out the next step.

OpenStudy (anonymous):

I have to take ECA's soon

OpenStudy (anonymous):

x-4=10. so what is x?

OpenStudy (anonymous):

14 as it appears to be.

OpenStudy (anonymous):

Correct 14. So... second please...

OpenStudy (anonymous):

wait I'm confused what did you do to the problem

OpenStudy (anonymous):

WAIT WAIT WAIT.... I used the wrong formula!!! Hold on and I will fix it.

OpenStudy (anonymous):

Sorry I am really tired... My brain is fused... Lol I am correcting it.

OpenStudy (anonymous):

SHOOT!!!! I have to finish my extra credit!! I have one hour and over 200 questions!!! Sorry dude, I forgot... I NEED to go. Re-bump your question, I am also asking someone from the chat group to help you.

jigglypuff314 (jigglypuff314):

2y=6x-4 divide both sides by 2 to solve for y what would you get?

OpenStudy (anonymous):

wait would I divide on the 6 or the 4

jigglypuff314 (jigglypuff314):

both :)

OpenStudy (anonymous):

so I would divide by 2 on all 3 of them

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (anonymous):

I got y=3x-2

jigglypuff314 (jigglypuff314):

that's right :) then plug in 3x - 2 for y into 3x-2y=10 so you get 3x - 2(3x-2) = 10 can you solve for x from there? :)

OpenStudy (anonymous):

would I multiply 3x-2 times 3x-2

jigglypuff314 (jigglypuff314):

distribute the negative 2 into (3x-2) so 3x - 2(3x-2) = 10 becomes 3x - 6x + 4 = 10 now add like terms

OpenStudy (anonymous):

I got 9x from adding 3x+6x

jigglypuff314 (jigglypuff314):

3x - 6x what is 3 minus 6?

OpenStudy (anonymous):

3-6 is negative 3

jigglypuff314 (jigglypuff314):

yes :) so -3x + 4 = 10 now subtract 4 from both sides :)

OpenStudy (anonymous):

-3x=6

jigglypuff314 (jigglypuff314):

good! now divide both sides by negative 3

OpenStudy (anonymous):

I would get x=-3

jigglypuff314 (jigglypuff314):

6 divided by (-3) = ?

OpenStudy (anonymous):

-2

jigglypuff314 (jigglypuff314):

yes! so x = -2 :) do you understand that?

OpenStudy (anonymous):

its somewhat comin to me

jigglypuff314 (jigglypuff314):

ok then to get y plug in x = -2 into the equation y = 3x - 2 so y = 3(-2) - 2 = ?